2. The mean height of Grade 11 students at a certain senior high school is 164 cm. and the standard deviation is 10 cm. Assuming that the heights are normally distributed, what percent of the height is

a. greater the 168 cm?

b. less than 150 cm

c. between 150 and 168


1
Expert's answer
2022-04-27T00:59:03-0400

We have a normal distribution, "\mu=164, \sigma=10."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}."


"\text{a. }z_1=\cfrac{168-164}{10}=0.40;
\\P(X>168)=P(Z>0.40)=\\
=1-P(Z<0.40)=\\
=1-0.6554=0.3446=34.46\%\text{ (from z-table)}."


"\text{b. }z_2=\cfrac{150-164}{10}=-1.40;
\\P(X<150)=P(Z<-1.40)=\\
=0.0808=8.08\%\text{ (from z-table)}."


"\text{c. }P(150<X<168)=\\
=P(-1.40<Z<0.40)=\\
=P(Z<0.40)-P(Z<-1.40)=\\
=0.6554-0.0808=0.5746=57.46\%."